# Text-Queried Audio Source Separation via Hierarchical Modeling

### v1

@@ V-D1 Global semantic feature evaluation
TABLE VI: Audio captioning on AudioCaps and Clotho
[…]
@@
AF 0.121 0.170 0.336
V-D2 Local semantic feature evaluation
We evaluate our pretrained AudioMAE model on the HEAR benchmark [67] and three tasks, ESC-50, SC-5h and NS-5h, where only the downstream model is trained. As shown in Table VII, our pretrained model achieves good performance, outperforming the official AudioMAE-B [45].
TABLE VII: Evaluation of AudioMAE on HEAR benchmark
[…]
AudioMAE-B [45] 57.6 33.9 61.4
@@ V-D2 Local semantic feature evaluation
V-E Neural codec performance
In Table VIII, we assess the reconstruction quality of general audio by our neural codec TF-Codec. We randomly sample 1000 audios from AudioSet validation set for evaluation. All models have a bitrate of 6 kbps. We can see that although our general audio TF-codec has a low latency by using a causal structure and much less parameters, it achieves good performance, superior than EnCodec [68]. DAC [69] achieves the best performance but it is much heavier with a non-causal structure. It should be noted that we utilize TF-Codec just as a proof of concept, and future works may utilize any non-causal codec with superior performances.
TABLE VIII: Neural codec performance on general audio at 6kbps
[…]
@@ VI Conclusion and future works
In this study, we propose a hierarchical modeling and separation framework for text-queried audio source separation, decoupling multi-level semantic feature separation and acoustic reconstruction. Leveraging Q-Audio for global-semantic modeling on top of AudioMAE for structure-preserving representations, our model achieves superior separation performance and semantic correctness, outperforming existing methods. The instruction parser enhances flexibility in handling diverse text queries with frozen LLMs. In future work, we will scale up the model in training data coverage which incorporates speech as well, and explore more fantastic editing options besides separation.

### v2

@@ V-F1 Global semantic feature evaluation
TABLE VII: Audio captioning on AudioCaps and Clotho
Model METEOR(){($\uparrow$)} SPICE(){($\uparrow$)} SPIDEr(){($\uparrow$)}
Audio Flamingo 0.121 0.170 0.336
@@
V-F2 Local semantic feature evaluation
We evaluate our pretrained AudioMAE model on the HEAR benchmark [turian2022hear] and three tasks, ESC-50, SC-5h and NS-5h, where only the downstream model is trained. As shown in Table VIII, our pretrained model achieves good performance, outperforming the official AudioMAE-B [huang2022masked]. This improvement can be attributed to two designs inspired by MW-MAE [yadav2023masked], i.e. employing a smaller decoder to encourage the encoder to capture richer contextual representations, and removing the dataset-level normalization.
TABLE VIII: Evaluation of AudioMAE on HEAR benchmark
Model ESC-50(){($\uparrow$)} SC-5h(){($\uparrow$)} NS-5h(){($\uparrow$)}
AudioMAE-B [huang2022masked] 57.6 33.9 61.4
@@ V-F2 Local semantic feature evaluation
V-G Neural codec performance
In Table IX, we assess the reconstruction quality of general audio by our neural codec TF-Codec. We randomly sample 1000 audios from AudioSet validation set for evaluation. All models have a bitrate of 6 kbps. We can see that although our general audio TF-codec has a low latency by using a causal structure and much less parameters, it achieves good performance, superior than EnCodec [defossez2022high]. DAC [kumar2023high] achieves the best performance but it is much heavier with a non-causal structure. It should be noted that we utilize TF-Codec just as a proof of concept, and future works may utilize any non-causal codec with superior performances.
TABLE IX: Neural codec performance on general audio at 6kbps
Model #param Causal MEL-D(){($\downarrow$)} LSD(){($\downarrow$)} VISQOL(){($\uparrow$)}
@@ V-G Neural codec performance
V-H Complexity
We compare the computational complexity with UNet-style separation method AudioSep and generative model based approach FlowSep in Table X. Here “# Param_sep” and “# Param_pretrained” denote the model parameters for separation module and other pretrained text or audio encoder/decoders, respectively. It is noted that although our HSM-TSS has more parameters than AudioSep, its complexity in number of FLOPs is much lower than both AudioSep and FlowSep as AudioSep uses a heavy ResUNet module for separation and FlowSep utilizes several steps, default to 20 in Table X, to generate the separated audio feature. We utilize fvcore [^4] https://github.com/facebookresearch/fvcore to calculate the FLOPs for each model.
TABLE X: Complexity comparisons for a 10-second audio
Model FLOPs/G # Param_sep/M # Param_pretrained/M
AudioSep 333.7 39.0 199.6
FlowSep 2046.9 265.6 410.5
HSM-TSS 183.8 129.9 378.3
@@ VI Conclusion and future works
In this study, we propose a hierarchical modeling and separation framework for text-queried audio source separation, decoupling multi-level semantic feature separation and acoustic reconstruction. Leveraging Q-Audio for global-semantic modeling on top of AudioMAE for structure-preserving representations, our model achieves superior separation performance and semantic correctness, outperforming existing methods. In future work, we will scale up the model in training data coverage, which incorporates speech as well, and explore a wider range of instruction-based editing capabilities beyond separation, such as style transfer or in-place modification.